# RIS-aided Wireless Communication with Movable Elements: Geometry Impact on Performance

### v1

@@ Abstract
Reconfigurable Intelligent Surfaces (RIS) are known as a promising technology to improve the performance of wireless communication networks, and have been extensively studied. Movable Antennas (MA) are a novel technology that fully exploits the antenna placement for enhancing the system performance. This article aims at evaluating the impact of transmit power and number of antenna elements on the outage probability performance of an MA-enabled RIS structure (MA-RIS), compared to existing Fixed-Position Antenna RIS (FPA-RIS). The change in geometry caused by the movement of antennas and its implications for the effective number of illuminated elements, are studied for 1D and 2D array structures. Our numerical results confirm the performance advantage provided by MA-RIS, achieving 24% improvement in outage probability, and 2 dB gain in Signal-to-Noise Ratio (SNR), as compared to FPA-RIS.
Index Terms:
RIS, Movable antenna, Fixed-position antenna, Performance analysis, Outage probability.
@@ I Introduction
A comprehensive channel model tailored for Movable Antenna (MA) systems, grounded in field-response principles, is introduced in [1]. Evaluations of MA system performance enhancements are conducted in [2], indicating promising outcomes compared to Fixed-Position Antennas (FPA) setups. These enhancements span various metrics such as channel capacity, sum-rate, array gain for beamforming in desired directions, and achievable rate [3], achieved through meticulous design and optimization of antenna positions. However, prevailing studies often assume uniform activity across all elements in RIS, overlooking the intricacies of geometry and element distribution.
@@
This simplification overlooks the dynamic nature of RIS operations, where not all elements are active simultaneously. Consequently, it prompts a novel avenue of exploration concerning the accurate identification of active elements, and determining the optimal number needed for peak performance in MA-RIS configurations. This arises from the anticipated deployment of RISs across expansive structures like buildings, where only portions of the RIS area would be illuminated at any given time. In [4], the notion of effective element count is introduced, tied to the illuminated area. The region illuminated by the main lobe of a transmitted beam is typically smaller than the entire MA-RIS area, often approximated as a conical or elliptical shape. Considering that highly directive beams facilitate communication between the Base Station (BS) and MA-RIS, at least 97% of the transmit energy typically resides within the Half-Power Beamwidth (HPBW) area of the main BS to MA-RIS beam [4].
@@
The efficiency of RIS is significantly influenced by the geometric arrangement of reflective elements and the compactness of the RIS panel in the near field, as indicated in [4], a notion further supported by [5] in the context of 3D RIS deployment. However, existing studies primarily focus on FPA-RIS structures, leaving a gap in understanding the impact of changes in the geometry of reflective element deployment in MA-RIS structures.
@@
An inherent challenge in adopting FPA-RIS structures lies in the potential phase distribution offset across various cascaded source-reflective element-destination channels, necessitating distinct patterns of optimal non-uniform discrete phase shifts for each element within the RIS. This phenomenon escalates manufacturing costs, particularly when dealing with a large number of RIS elements [6]. Conversely, fewer antennas require less physical space for installation, rendering the system more compact and potentially suitable for space-constrained applications. Additionally, the number of elements within an RIS panel significantly influences its output performance. For instance, [7] examines how the number of elements impacts system sum rate, aiming to identify optimal configurations. Results indicate a positive correlation between the number of reflecting elements and system sum-rate enhancement.
@@
One such potential solution involves integrating MA within the RIS panel, which may achieve comparable or superior system performance while utilizing fewer reflecting elements compared to existing FPA-RIS configurations.
Driven by the aforementioned considerations, our investigation delves into understanding how the number of elements and transmit power influence outage probability and SNR performance in MA-RIS systems. We conduct a comprehensive comparative assessment elucidating the impacts of MA element count and transmit power on system robustness, with a specific focus on outage probability as a key performance indicator. The results show that the MA-RIS could achieve lower outage probability than FPA-RIS with less antenna numbers in both 1D and 2D configurations.
Driven by the aforementioned considerations, our investigation delves into understanding how the number of elements and transmit power influence outage probability and SNR performance in MA-RIS systems, focusing on sub-6-GHz array antennas in near-field scenarios. The primary contribution of this paper is two-fold. Firstly, we provide a theoretical framework for delineating the illuminated area and determining the effective number of reflecting antenna elements across different MA-RIS configurations (1D versus 2D), factoring in geometric variables such as distance, HPBW, and angle of arrival and departure. Secondly, we integrate environmental blockage effects into our analysis of LoS channels, enabling the derivation of outage probability metrics for varying effective numbers of MA-RIS elements. We conduct a comprehensive comparative assessment elucidating the impacts of MA element count and transmit power on system robustness, with a specific focus on outage probability as a key performance indicator.
@@ II-A Channel model
In a downlink scenario leveraging MA-RIS, we explore a system where $K$ User Equipments (UE) receive data via reflections from a RIS panel composed of $N$ movable antenna elements as shown in Fig. 1. The BS is positioned at $\textbf{b}_{0}=[x_{b},y_{b},z_{b}]^{T}$, and equipped with $M$ FPAs transmitting signal towards MA-RIS. The position of the $m^{th}$ FPA at BS can be represented as $\textbf{b}_{m}=[x^{b}_{m},y^{b}_{m},z^{b}_{m}]^{T}$, for $m\in\mathcal{M}\triangleq\left\{1,\dots,M\right\}$. As for UEs, denoted by $k\in\mathcal{K}=\{1,\dots,K\}$, each is equipped with a single FPA. The position of the $k^{th}$ UE is denoted as $\textbf{u}_{k}=\left[x^{u}_{k},y^{u}_{k},z^{u}_{k}\right]^{T}$. The MA-RIS panel, featuring with a certain length, enables 1D and 2D alignment MA moving horizontally from the left panel edge to the right panel edge. The 1D and 2D structures are equipped with $1\times N$ and $N_{x}\times N_{y}$ MAs separately, where we let $N=N_{x}\times N_{y}$ to make sure the 1D and 2D structured MA-RIS have the same number of antennas. The position of the MA-RIS center is denoted as $\textbf{r}_{c}=\left[x_{r},y_{r},z_{r}\right]^{T}$. The position of the $n^{th}$ MA is represented by Cartesian coordinates $\textbf{r}_{n}=\left[x^{r}_{n},y^{r}_{n},z^{r}_{n}\right]^{T}\in C_{r}$ for $n\in\mathcal{N}\triangleq\left\{1,\dots,N\right\}$, where $C_{r}$ denotes the 3D region within which the MAs can move freely [10, 1].
Figure 1: Illustration of the system model. In the MA-RIS panel, yellow blocks depict all the $N$ MA units moving with speed $v$, while the green blocks signify the effective antennas. The solid blue ellipse is an enlarged representation of the hollow ellipse which represents the illumination area labeled as $S_{m}$, with semi-major axis $a$ and semi-minor axis $b$, and represents the illumination area, showing the HPBW $\phi_{0}$ from the BS. The terms $\varphi_{n}^{r}$ and $\varphi_{n}^{t}$ denote the AoA and AoD associated with each antenna element, respectively.
@@
Given the dominant LoS link with a signal path between both the BS to MA-RIS and MA-RIS to UE channels, the normalized wave vector of one single transmit path can be represented by $\textbf{n}=[\mathrm{sin}\theta_{b}\mathrm{cos}\varphi_{b},\mathrm{cos}\theta_{b},\mathrm{sin}\theta_{b}\mathrm{sin}\varphi_{b}]^{T}$, where $\varphi_{b}$ and $\theta_{b}$ represent the Elevation of Departure (EoD) and Azimuth of Departure (AoD) from the BS to MA-RIS link, respectively. Therefore, following [11], the path difference relative to the reference point $\textbf{b}_{0}=[0,0,0]^{T}$ is denoted as $\rho(\textbf{b}_{m})=\textbf{n}^{T}(\textbf{b}_{m}-\textbf{b}_{0})=x^{b}_{m}\mathrm{sin}\theta_{b}\mathrm{cos}\varphi_{b}+y^{b}_{m}\mathrm{cos}\theta_{b}+z^{b}_{m}\mathrm{sin}\theta_{b}\mathrm{sin}\varphi_{b}$. Accordingly, the Field-Response Vector (FRV) at the BS, considering the linear FPA array, is expressed as:
$\textbf{s}(\mathbf{b}_{m})=\left[e^{j\frac{2\pi}{\lambda}\rho(\textbf{b}_{1})},e^{j\frac{2\pi}{\lambda}\rho(\textbf{b}_{2})},\cdots,e^{j\frac{2\pi}{\lambda}\rho(\textbf{b}_{M})}\right]^{T}$ (1)
where $\lambda$ is the carrier length. Considering the 2D plane deployment of MA-RIS [4], therefore, the FRV for the transmit path from the MA-RIS at position $\textbf{r}_{n}$, is simplified to a scalar value in the context of LoS channel model [12], yielding:
$g^{r}(\textbf{r}_{n})=e^{j\frac{2\pi}{\lambda}x^{r}_{n}\mathrm{cos}\varphi_{n}^{r}}$ (2)
where $\varphi_{n}^{r}$ represents the Azimuth of Arrival (AoA) for each MA-RIS element. By stacking together the $g^{r}(\textbf{r}_{n})$’s of all $N$ transmit MA, we obtain the collective FRV, is given by
$\textbf{G}^{r}(\textbf{r})\triangleq\left[g^{r}(\textbf{r}_{1}),g^{r}(\textbf{r}_{2}),\dots,g^{r}(\textbf{r}_{N})\right]^{T}\in\mathbb{C}^{N\times 1}$ (3)
Similarly, by defining $g^{t}(\textbf{r}_{n})=e^{j\frac{2\pi}{\lambda}x^{r}_{n}\mathrm{cos}\varphi_{n}^{t}}$, the FRV of MA-RIS with $\varphi_{n}^{t}$ representing the AoD from each MA-RIS element to UE link is
$\textbf{G}^{t}(\textbf{r})\triangleq\left[g^{t}(\textbf{r}_{1}),g^{t}(\textbf{r}_{2}),\dots,g^{t}(\textbf{r}_{N})\right]^{T}\in\mathbb{C}^{N\times 1}$ (4)
The $k^{th}$ single UE is equipped with a single FPA with $\varphi_{k}^{u}$ representing the AoA to UE, and the corresponding FRV can be defined as
$f(\textbf{u}_{k})=e^{j\frac{2\pi}{\lambda}x^{u}_{k}\mathrm{cos}\varphi_{k}^{u}}$ (5)
The channel vector for the BS-RIS link can then be written as
$\mathbf{\rm h}_{br}=\textbf{G}^{r}(\textbf{r})^{H}\textbf{s}(\mathbf{b}_{m})$ (6)
while the channel vector for the RIS-UE link can be written as
$\mathbf{\rm h}_{ru}=f(\textbf{u}_{k})\textbf{G}^{t}(\textbf{r})$ (7)
As in [13, 6], we denote $\mathbf{\rm h}_{br}=\left[{h}_{br,1},\cdots,{h}_{br,N}\right]^{T}$ and $\mathbf{\rm h}_{ru}=\left[{h}_{ru,1},\cdots,{h}_{ru,N}\right]$, and elements in those channel vectors are separately given as
$h_{br,n}=\alpha_{n}e^{-j\frac{2\pi}{\lambda}\left(\rho(\textbf{b}_{m})+x^{r}_{n}\mathrm{cos}\varphi_{n}^{r}\right)}$ (8)
$h_{ru,n}=\beta_{n}e^{-j\frac{2\pi}{\lambda}\left(x^{r}_{n}\mathrm{cos}\varphi_{n}^{t}+x^{u}_{k}\mathrm{cos}\varphi_{k}^{u}\right)}$ (9)
where $\alpha_{n}$ and $\beta_{n}$ represent the channel’s response. Therefore, the received signal at UE $k$ can be expressed as follows:
$y_{k}=\sqrt{P_{r}}h_{ru,n}\boldsymbol{\Theta}h_{br,n}s+n_{0}$ (10)
where $P_{r}$ is the received power, $s\sim\mathcal{CN}(0,1)$ is the signal transmitted by the BS, $n_{0}$ is the additive complex Gaussian noise with zero mean and variance of $\sigma_{n}^{2}$, and $\boldsymbol{\Theta}=diag\left\{e^{j\theta_{1}},\cdots,e^{j\theta_{N}}\right\}$ is the phase shift matrix of MA-RIS. Letting $\varpi=\rho(\textbf{b}_{m})+x^{r}_{n}\mathrm{cos}\varphi_{n}^{r}+x^{r}_{n}\mathrm{cos}\varphi_{n}^{t}+x^{u}_{k}\mathrm{cos}\varphi_{k}^{u}$, $y_{k}$ can be expanded as
$y_{k}$ $=$ $\sqrt{P_{r}}h_{ru,n}\boldsymbol{\Theta}h_{br,n}s+n_{0}$
$=$ ${\sqrt{P_{r}}}\sum_{i=1}^{N_{eff}}\alpha_{i}\beta_{i}e^{j\left(\theta_{i}-\frac{2\pi}{\lambda}\varpi\right)}s+n_{0}$
where $N_{eff}$ is the effective number of illuminated elements determined by the illumination area in MA-RIS, leading to the UE’s received instantaneous SNR of
$\gamma={\frac{P_{r}}{\sigma_{n}^{2}}}\left|\sum_{i=1}^{N_{eff}}\alpha_{i}\beta_{i}e^{j\left(\theta_{i}-\frac{2\pi}{\lambda}\varpi\right)}\right|^{2}$ (12)
By inspecting Eq. (12), we note that $\rho$ reaches its maximum value in correspondence of:
@@
$\theta_{i}=\frac{2\pi}{\lambda}\left(\rho(\textbf{b}_{m})+x^{r}_{n}\mathrm{cos}\varphi_{n}^{r}+x^{r}_{n}\mathrm{cos}\varphi_{n}^{t}+x^{u}_{k}\mathrm{cos}\varphi_{k}^{u}\right)$ (13)
@@
Hence, we can get the maximum SNR as $\gamma_{max}=A^{2}\bar{\gamma}$, where $A=\sum_{i=1}^{N_{eff}}\alpha_{i}\beta_{i}$, and the average SNR is $\bar{\gamma}={\frac{P_{r}}{\sigma_{n}^{2}}}$.

### v2

@@ Abstract
Reconfigurable Intelligent Surfaces (RIS) are known as a promising technology that can improve the performance of wireless communication networks and have been extensively studied. Movable Antennas (MA) are a novel technology that fully exploits the antenna placement to enhance the system performance. This article aims to evaluate the impact of transmit power and number of antenna elements on the outage probability performance of an MA-enabled RIS structure (MA-RIS), compared to existing Fixed-Position Antenna RIS (FPA-RIS). The change in geometry caused by the movement of antennas and its implications for the effective number of illuminated elements, are studied for 1D and 2D array structures. Our numerical results confirm the performance advantage provided by MA-RIS, achieving 24% improvement in outage probability, and 2 dB gain in Signal-to-Noise Ratio (SNR), as compared to FPA-RIS.
@@ I Introduction
A comprehensive channel model tailored for Movable Antenna (MA) systems, grounded in field-response principles, is introduced in [18]. Evaluations of MA system performance enhancements are conducted in [17], indicating promising outcomes compared to Fixed-Position Antennas (FPA) setups. These enhancements span various metrics such as channel capacity, sum-rate, array gain for beamforming in desired directions, and achievable rate [7], achieved through meticulous design and optimization of antenna positions. However, prevailing studies often assume uniform activity across all elements in RIS and mainly consider the MA is deployed at transmitter/receiver/relay, overlooking the intricacies of geometry and element distribution [9].
@@
This simplification overlooks the dynamic nature of RIS operations, where not all elements are active simultaneously. Consequently, it prompts a novel avenue of exploration concerning the accurate identification of active elements, and determining the optimal number needed for peak performance in MA-RIS configurations. This arises from the anticipated deployment of RISs across expansive structures like buildings, where only portions of the RIS area would be illuminated at any given time. In [6], the notion of effective element count is introduced, tied to the illuminated area. The region illuminated by the main lobe of a transmitted beam is typically smaller than the entire MA-RIS area, often approximated as a conical or elliptical shape. Considering that highly directive beams facilitate communication between the Transmitter (Tx) and MA-RIS, at least 97% of the transmit energy typically resides within the Half-Power Beamwidth (HPBW) area of the main Tx to MA-RIS beam [6].
@@
The efficiency of RIS is significantly influenced by the geometric arrangement of reflective elements and the compactness of the RIS panel in the near field, as indicated in [6], a notion further supported by [4] in the context of 3D RIS deployment. Optimization efforts are carried out in RIS systems equipped with movable antennas [12]. However, existing studies primarily focus on FPA-RIS structures, leaving a gap in understanding the impact of changes in the geometry of reflective element deployments in MA-RIS structures.
@@
An inherent challenge in adopting FPA-RIS structures lies in the potential phase distribution offset across various cascaded source-reflective element-destination channels, necessitating distinct patterns of optimal non-uniform discrete phase shifts for each element within the RIS. This phenomenon escalates manufacturing costs, particularly when dealing with a large number of RIS elements [5]. Conversely, fewer antennas require less physical space for installation, rendering the system more compact and potentially suitable for space-constrained applications. Additionally, the number of elements within an RIS panel significantly influences its output performance. For instance, [16] examines how the number of elements impacts the system sum rate, aiming to identify optimal configurations. Results indicate a positive correlation between the number of reflecting elements and system sum-rate enhancement.
@@
One such potential solution involves integrating MA within the RIS panel, which may achieve comparable or superior system performance while utilizing fewer reflecting elements compared to existing FPA-RIS configurations. This could lead to several advantages: a) Fewer Reflecting Elements: Metamaterial antennas can often achieve the same or better performance compared to conventional antennas while using fewer components. This means an MA-integrated RIS panel could potentially use fewer reflecting elements than a standard RIS panel, leading to a more compact and efficient design. b) Comparable or Superior System Performance: The unique properties of metamaterials might enable the MA-integrated RIS panel to manipulate radio waves in more sophisticated ways, potentially improving the overall performance of the RIS system in terms of signal strength, beamforming capabilities, or interference reduction.
Driven by the aforementioned considerations, our investigation delves into understanding how the number of elements and transmit power influence outage probability and SNR performance in MA-RIS systems, focusing on sub-6-GHz array antennas in near-field scenarios. The primary contribution of this paper is two-fold. Firstly, we provide a theoretical framework for delineating the illuminated area and determining the effective number of reflecting antenna elements across different MA-RIS configurations (1D versus 2D), factoring in geometric variables such as distance, HPBW, and angle of arrival and departure. Secondly, we integrate environmental blockage effects into our analysis of LoS channels, enabling the derivation of outage probability metrics for varying effective numbers of MA-RIS elements. We conduct a comprehensive comparative assessment elucidating the impacts of MA element count and transmit power on system robustness, with a specific focus on outage probability as a key performance indicator. The results show that the MA-RIS could achieve a lower outage probability than FPA-RIS with a lower number of antennas in both 1D and 2D configurations.
@@ II-A Channel model
In this section, we introduce the system model and channel model. The system model depicted in Fig. 1, we model the partial RIS illumination as an elliptical area, which is essential for determining the effective number of elements [10].
The MA-RIS panel, featuring with a certain length, enables 1D and 2D alignment MAs moving horizontally from the left panel edge to the right panel edge. The 1D and 2D structures are equipped with $1\times N$ and $N_{x}\times N_{y}$ MAs separately, where we let $N=N_{x}\times N_{y}$ to make sure the 1D and 2D structured MA-RIS have the same number of MAs.
@@
We consider the simplified LoS scenario with only one transmit and receive path between Tx and RIS, and the same for the RIS to Rx path. The position of Tx can be represented as $\textbf{b}=[x_{b},y_{b},0]^{T}$. Thus, the normalized wave vector of one single transmit path can be represented by $\textbf{n}=[\mathrm{sin}\theta_{b}\mathrm{cos}\varphi_{b},\mathrm{cos}\theta_{b},\mathrm{sin}\theta_{b}\mathrm{sin}\varphi_{b}]^{T}$, where $\theta_{b}$ and $\varphi_{b}$ represent the Elevation of Departure (EoD) and Azimuth of Departure (AoD) from the Tx to MA-RIS link, respectively. Therefore, following [8], the relative path difference from b to reference point $\textbf{b}_{0}=[0,0,0]^{T}$ is denoted as $\rho(\textbf{b})=\textbf{n}^{T}(\textbf{b}-\textbf{b}_{0})=x_{b}\mathrm{sin}\theta_{b}\mathrm{cos}\varphi_{b}+y_{b}\mathrm{cos}\theta_{b}$. The resulting phase difference is $\frac{2\pi}{\lambda}\rho(\textbf{b})$. Moreover, the array steering vector at Tx is defined as follows
@@
$s(\textbf{b})=e^{j\frac{2\pi}{\lambda}\rho(\textbf{b})}=e^{j\frac{2\pi}{\lambda}\left(x_{b}\mathrm{sin}\theta_{b}\mathrm{cos}\varphi_{b}+y_{u}\mathrm{cos}\theta_{b}\right)}\,$, (1)
where $\lambda$ is the carrier wavelength.
The center position of the MA-RIS panel is denoted as $\textbf{r}_{c}=\left[x_{r},y_{r},z_{r}\right]^{T}$. The position of the $n$-th MA is represented as $\textbf{r}_{n}=\left[x^{r}_{n},y^{r}_{n},z^{r}_{n}\right]^{T}\in C_{r}$ for $n\in\mathcal{N}\triangleq\left\{1,\dots,N\right\}$, where $C_{r}$ denotes the region within which the MAs can move freely [1, 18]. Therefore, we get the path difference from Tx to the $n$-th MA as $g^{r}(\textbf{r}_{n})=e^{j\frac{2\pi}{\lambda}\rho^{r}(\textbf{r}_{n})}=e^{j\frac{2\pi}{\lambda}\left(x^{r}_{n}\mathrm{sin}\theta_{n}^{r}\mathrm{cos}\varphi_{n}^{r}+y^{r}_{n}\mathrm{cos}\theta_{n}^{r}\right)}$, where $\theta^{r}_{n}$ and $\varphi_{n}^{r}$ respectively represents the Elevation of Arrival (EoA) and Azimuth of Arrival (AoA) for each MA-RIS element, and $\rho^{r}(\textbf{r}_{n})$ is the path difference from $\textbf{r}_{n}$ to reference point relative to the received path.
Figure 1: Illustration of the system model. In the MA-RIS panel, yellow blocks depict all the $N$ MA units moving with speed $v$, while the green blocks signify the effective antennas. The solid blue ellipse is an enlarged representation of the hollow ellipse which represents the illumination area labeled as $S_{m}$, with semi-major axis $a$ and semi-minor axis $b$, and represents the illumination area, showing the HPBW $\phi_{0}$ from the Tx.
By stacking together the $g^{r}(\textbf{r}_{n})$ of all $N$ MAs, we obtain the collective array response vector of the receive MA is defined as follows
$\textbf{G}^{r}(\textbf{r})\triangleq\left[g^{r}(\textbf{r}_{1}),g^{r}(\textbf{r}_{2}),\dots,g^{r}(\textbf{r}_{N})\right]\,.$ (2)
Similarly, by defining $g^{t}(\textbf{r}_{n})=e^{j\frac{2\pi}{\lambda}\rho^{t}(\textbf{r}_{n})}=e^{j\frac{2\pi}{\lambda}\left(x^{r}_{n}\mathrm{sin}\theta_{n}^{t}\mathrm{cos}\varphi_{n}^{t}+y^{r}_{n}\mathrm{cos}\theta_{n}^{t}\right)}$, and $\rho^{t}(\textbf{r}_{n})$ is the path difference from $\textbf{r}_{n}$ to reference point relative to the transmitted path. The array response of MA-RIS with $\theta^{t}_{n}$ and $\varphi_{n}^{t}$ representing the EoD and AoD from $n$-th MA-RIS element to Rx link is defined as
$\textbf{G}^{t}(\textbf{r})\triangleq\left[g^{t}(\textbf{r}_{1}),g^{t}(\textbf{r}_{2}),\dots,g^{t}(\textbf{r}_{N})\right]\,.$ (3)
The Rx is fixed at $\textbf{u}=\left[x_{u},y_{u},z_{u}\right]^{T}$ and $\theta^{u}$ and $\varphi^{u}$ represent the EoA and AoA to Rx. Then with the path difference $\rho^{t}(\textbf{u})$, the corresponding FRV relative to the reference point is expressed as
$f(\textbf{u})=e^{j\frac{2\pi}{\lambda}\rho^{t}(\textbf{u})}=e^{j\frac{2\pi}{\lambda}\left(x_{u}\mathrm{sin}\theta^{u}\mathrm{cos}\varphi^{u}+y_{u}\mathrm{cos}\theta^{u}\right)}\,.$ (4)
The small-scale fading channel vector for the Tx to MA-RIS link can then be written as
$\rm{\mathbf{h}}_{br}=\alpha_{n}\mathbf{G}^{r}(\mathbf{r})^{H}s(\mathbf{b})\,$, (5)
while the channel vector for the MA-RIS to Rx link can be written as
$\rm{\mathbf{h}}_{ru}=\beta_{n}f(\mathbf{u})\mathbf{G}^{t}(\mathbf{r})\,$, (6)
where $\alpha_{n}$ and $\beta_{n}$ represent the channel’s response coefficients.
Denoted by ${\rm s}\in\mathbb{C}^{M\times 1}$ the input of the downlink transmission, then the received signal at Rx is expressed as
$y_{k}=\sqrt{P_{r}}{\rm{{\mathbf{h}}_{ru}\boldsymbol{\Theta}\rm{\mathbf{h}}_{br}s}}+n_{0}\,$, (7)
where $P_{r}$ is the received power, $n_{0}$ is the additive complex Gaussian noise with zero mean and variance of $\sigma_{n}^{2}$, and $\boldsymbol{\Theta}=\mathrm{diag}\left\{e^{j\theta_{1}},\cdots,e^{j\theta_{N}}\right\}$ is the phase shift matrix of MA-RIS. Letting $\varpi=\rho(\textbf{b})+\rho^{r}(\textbf{r}_{n})+\rho^{t}(\textbf{r}_{n})+\rho^{t}(\textbf{u})$, following the SNR maximization rule in [2], $y_{k}$ can be expanded as
$y_{k}$ $=\sqrt{P_{r}}{\rm{{\mathbf{h}}_{ru}\boldsymbol{\Theta}\rm{\mathbf{h}}_{br}s}}+n_{0}$
$={\sqrt{P_{r}}}\sum_{n=1}^{N_{eff}}\alpha_{n}\beta_{n}e^{j\left(\theta_{i}-\frac{2\pi}{\lambda}\varpi\right)}s+n_{0}\,$,
where $N_{eff}$ is the effective number of elements determined by the illumination area in MA-RIS, leading to the Rx’s received instantaneous SNR as
$\gamma={\frac{P_{r}}{\sigma_{n}^{2}}}\left|\sum_{n=1}^{N_{eff}}\alpha_{n}\beta_{n}e^{j\left(\theta_{i}-\frac{2\pi}{\lambda}\varpi\right)}\right|^{2}\,.$ (9)
By inspecting (9), we note that $\gamma$ reaches its maximum value in correspondence of
$\theta_{i}=\frac{2\pi}{\lambda}\left(\rho(\textbf{b})+\rho^{r}(\textbf{r}_{n})+\rho^{t}(\textbf{r}_{n})+\rho^{t}(\textbf{u})\right)\,.$ (10)
@@
Hence, we can get the maximum SNR as $\gamma_{max}=A^{2}\bar{\gamma}$, where $A=\sum_{n=1}^{N_{eff}}\alpha_{n}\beta_{n}$, and the average SNR is $\bar{\gamma}={\frac{P_{r}}{\sigma_{n}^{2}}}$.